# Supplementary Information

### v1

@@ Results
Simulated datasets were generated based on two-state and four-state fluorescence lifetime–intensity models to evaluate the performance of the clustering algorithm. For all sets of simulated data, the background level was chosen to be non-zero to mimic actual experiments where sources of external noise are inevitable. To benchmark the method under ideal conditions, normally distributed noise was added around the mean values of each state, constrained by their respective SDs, and these were used as the initial data from which the intensities and lifetimes were sampled. This normally distributed noise mimics the dense populations of SMS data. Specifically, for the lower intensity state, the mean and SD used for sampling of the intensity were initially set to $500\,\text{counts/s}$ (cps) and $50\,\text{cps}$, respectively, and for the higher state they were initially set to $2500\,\text{cps}$ and $150\,\text{cps}$, respectively. Under these conditions, the algorithm is expected to produce centrally clustered results that accurately reflect the underlying model structure. Figure 1 displays the results for a system of $300$ simulated particles, each having a 10-minute trace. Figure 1A shows an example of a simulated intensity trace, displaying typical photoblinking behavior, i.e., abrupt transitions between an “on" and an “off" state, where the off-state corresponds to the background noise intensity. The lifetime–intensity distribution from a set of 300 simulated traces is shown in Figure 1B, clearly revealing the two populations. The clustering protocol was applied to this set of lifetime–intensity traces, and the corresponding BIC plot is shown in Figure 1C, indicating that two clusters are identified as the statistically most favorable outcome. Figure 1D shows the optimally-clustered cluster centers in the center of each population and the 0.9 confidence ellipses using the covariance and mean for each cluster component, providing a representative illustration of the data points associated with each cluster.
@@
Figure 1: Results of the clustering protocol performed on the two-state simulated dataset with 300 simulated particles, each with a time trace of 10 min. A. Representative example of a simulated two-state intensity trace (using $40$-ms binning for display). B. Lifetime–intensity distribution of this set of 300 simulated particles. C. Corresponding BIC score plot. D. Outcome of clustering, showing cluster centers (black crosses) and 0.9 confidence ellipses, estimated from the mean and covariance.
@@
To incorporate the effect of the excitation light polarization and the particle orientation, we applied Eq. (2) to the different intensity states from Figure 1 for circularly polarized light ($\epsilon=1$) and a set of randomized dipole orientations. The excitation probability is dictated only by the polar angle $\theta$ in the current case (for circularly polarized light), following a $\sin^{2}$ dependence on $\theta$, i.e., the dipole orientation’s projection to the focal plane. This produces a strong broadening of the population distribution toward lower fluorescence intensities. An example intensity trace, for a particle assumed to have a fixed orientation defined by $\theta=\frac{2\pi}{3}$ and $\phi=\frac{3\pi}{4}$, yielding an excitation probability of $P_{\text{{Exc}}}=0.75$ according to Eq. (2), is shown in Figure 2A where photoblinking occurs between the off-state and the reduced on-state corresponding to a simulated maximum emission intensity of 1875 cps. The smeared lifetime–intensity distribution for a system of 300 randomly oriented particles is shown in Figure 2B, where significant broadening of the unquenched state to lower intensities, down to the background noise level, is visible. The average of all the background-subtracted intensities of the unquenched population is 830 cps, which is in agreement with the expected 1/3 statistical average of the maximum intensity of 2500 cps.
Applying the clustering protocol to this smeared dataset gives the BIC score plot in Figure 2C. The elbow points, corresponding to the points of greatest gradient increase, are at 2 and 3. These points are considered to be the first major local minima, indicating that either two or three clusters are the optimal number. The result for two clusters is shown in Figure 2D, where the cluster of the higher state is centered at its mean intensity, and that of the lower state coincides with the densely populated off-state. The ellipses around cluster centers assign the data points to the two respective clusters at a confidence level of $0.9$, providing visual insight into the cluster structure. The result for three clusters is shown in Figure 2E. The long-lifetime data are now split into two clusters, with the first cluster center (cluster 2) around the dense population at the background intensity, and the second cluster center (cluster 3) at the mean intensity of the remaining data. Visual inspection suggests that cluster 3 should rather be aligned with the population at 2500 cps. However, since the density of data at 2500 cps is significantly lower than at 500 cps, the algorithm assigns a cluster center in between the two populations, slightly closer to the higher-density population.
@@
Figure 2: Influence of the excitation light polarization on the simulated intensity of a molecular dipole. A. Example intensity trace (with 40 ms binning) for circularly polarized light and a dipole orientation of $\theta=\frac{2\pi}{3}$ and $\phi=\frac{3\pi}{4}$, simulated for a time of 10 min. B. Lifetime–intensity distribution of a set of 300 fixed but randomly oriented dipoles. C. Corresponding BIC score plot. D. Cluster centers for a two-cluster model. E. Excitation probability histogram of a set of $10^{6}$ randomly oriented dipoles for circularly ($\epsilon=1$, red), elliptically ($\epsilon=0.7$, blue), and linearly ($\epsilon=0$, green) polarized light. F. Clustering results using three clusters, with cluster centers indicated by black crosses. Ellipses indicate the region of data that belongs to each cluster, at a confidence of $0.9$.
To evaluate the clustering protocol on a more complicated system, the protocol was applied to a simulated four-state system that was generated by adding two additional states, with intensities of 1200 cps and 1800 cps and corresponding SDs of 120 cps and 180 cps, respectively, to the original two-state dataset. The lifetime–intensity distribution of this dataset shows four well-defined states (Fig. S1A), which are seamlessly resolved by our clustering protocol (Fig. S1B–C). It is clear that the clustering protocol identified cluster centers corresponding exactly to the centers of the dense data regions.
Our benchmarking of the clustering protocol on simulated datasets was successful: the well-defined two- and four-state models were clustered exceptionally well with cluster centers coinciding with dense data regions. For the smeared dataset, the protocol correctly assigned cluster centers based on the underlying density of the data distribution, thus identifying the statistically most meaningful populations, despite one of the clusters appearing visually counterintuitive. In addition, the confidence ellipses not only assign the data to the most appropriate clusters but also depict the spread in the data connected to each cluster.
@@
We applied our clustering protocol to both the ungrouped and grouped-level Alexa data to demonstrate the effectiveness of the grouping technique described in Ref. (botha2024advanced). The Alexa dataset was chosen as a case study of rapid photoblinking, with dwell times in intensity states generally too short to be resolved visually or with a statistical change-point analysis (Fig. 3A). This was attained by deliberately not removing oxygen from the buffer solution. In addition, spin-coating of the sample introduced a random distribution of dipole orientations. We used circularly polarized excitation light to consider only the polar angle dependence of the dipole orientation, similarly to Figure 2. The resulting erratic intensity behavior and data smearing are shown in Figure 3, with the left column representing the resolved-level (ungrouped) data and the right column representing the grouped-level data. Clearly, change-point analysis of the data had difficulty resolving the actual intensity levels accurately (Fig. 3A, green trace), whereas this improved markedly after grouping the data using the built-in grouping feature of Full SMS (Fig. 3B, green trace). The lifetime–intensity distribution of the ungrouped data (Fig. 3C) was also significantly more spread out than that of the grouped data (Fig. 3D) and showed little discernible trend, making it difficult to distinguish between individual populations. In contrast, the unbiased grouping (Fig. 3D) statistically assigns similar resolved levels to a single grouped level, which effectively denoises the data and facilitates easier and more accurate clustering. Furthermore, the grouping also removed most of the short-lifetime data in the ungrouped distribution, which highlights the large fitting uncertainties when dealing with low photon budgets.
For both the ungrouped and grouped data, the BIC score plots indicate that three and five clusters correspond to the statistically most optimal values (Figs. 3E and F). The BIC score plot typically does not obey a monotonically decreasing trend and, therefore, could possess multiple local minima. Motivated by Occam’s razor, we adopt the first elbow points as the optimal solutions, which occur after the first steepest gradient descent in Figure 3E and at the first major local minimum in Figure 3F. This ensures the simplest model is selected to sufficiently describe the data without overfitting.
The resulting three-cluster analyses are shown in Figures 3G and H for the ungrouped and grouped data, respectively. The cluster centers of the ungrouped data (Fig. 3G) do not coincide with the densest populations due to the significant spread in the data; instead, the cluster centers correspond to the mean values of the data contained in the confidence ellipses. In contrast, the cluster centers of the grouped-level data correspond well to their visually expected positions (Fig. 3H). Clearly, grouping of the data leads to more streamlined and accurate clustering.
The choice of the first local minimum in the BIC analysis prevents overfitting. If the second minimum (corresponding to five clusters) is used instead (Fig. S2), state 2 in Figure 3H appears to be partitioned into multiple separate clusters, which is less likely to accurately describe the system. This inaccuracy is further highlighted by the considerably increased overlap between the confidence ellipses, creating significant ambiguity in cluster assignment.
@@
Figure 3: Comparison of the performance of the clustering protocol on the ungrouped (left column) and grouped (right column) Alexa data with 52 particles. A and B. Example intensity traces over 10-s windows, with grey representing the 40-ms binned intensity data and green showing the ungrouped (A) or grouped (B) trace. C and D. Lifetime–intensity distributions for the corresponding datasets. E and F. BIC score plots, indicating local minima used to determine the optimal number of clusters. G and H. Clustering outcomes using three clusters, with cluster centers indicated by black crosses, with corresponding 0.85 confidence ellipses.

### v2

@@ Results
Simulated datasets were generated from fluorescence lifetime–intensity models with varying numbers of underlying states to evaluate the clustering protocol under controlled conditions, with all analyses performed at a confidence level of $0.95$. Background noise was included in all datasets, and on- and off-state dwell times were drawn from exponential distributions. Alternative dwell-time models are presented in the SI. Normally distributed noise around each state’s mean, constrained by its standard deviation, was added to mimic dense SMS populations. To systematically assess the protocol’s limits, we started with simple, well-defined systems comprising equally populated, tight clusters (i.e., small sampling SDs) and progressively introduced additional complexity. This approach allows identification of the approximate point at which increasing overlap and diffuse populations compromise cluster recovery, thereby defining the “breaking point” of the algorithm. At selected points in this analysis, clustering performance was additionally compared against DPGMM, in which the effective number of components was inferred within a Bayesian framework, and K-means clustering, with the number of clusters determined using the elbow criterion.
Well-defined system with equal probability states
To simulate a two-state system (Fig. 1), intensities were sampled at $500\pm 120$ cps and $2600\pm 200$ cps, where the values denote the mean $\pm$ SD from normal distributions, while lifetimes were correspondingly sampled at $0.8\pm 0.22$ ns and $2.1\pm 0.3$ ns, respectively. For a four-state system (Fig. S1), two additional states were introduced with intensities of $1200\pm 240$ cps and $3100\pm 200$ cps, and corresponding lifetimes of $2.4\pm 0.2$ ns and $1.1\pm 0.2$ ns, respectively. Under these conditions, the clustering protocol is expected to recover well-defined populations that match the underlying model. Figure 1A shows a representative simulated intensity trace exhibiting rapid photoblinking, where the off-state corresponds to the background noise intensity. The lifetime–intensity distribution (Fig. 1B) clearly shows two populations. The IC score plots (AIC, BIC, and ICL; Fig. 1C) exhibit a sharp decrease to two clusters, with negligible improvement beyond these numbers. ICL deviates significantly from AIC and BIC beyond two clusters due to its additional penalty for cluster overlap. Cluster tightness, defined by the mean Mahalanobis distance, matched the expected value of $\langle D^{2}\rangle=2$, and the average fraction of points outside the $0.95$ confidence ellipse was FO =0.05=0.05, confirming that the clusters are compact and accurately reflect the underlying populations. Figure 1D shows the optimally clustered two-state system, with cluster centers at $(500\text{ cps},0.8\text{ ns})$ and $(2601\text{ cps},2.1\text{ ns})$, and ARI =1=1, indicating perfect ground-truth recovery. The clustering of the additional well-defined four-state system, which yielded identical results ($\langle D^{2}\rangle=2$, FO = 0.05, ARI =1=1), shown in Figure S1, indicates that higher-order well-defined states do not affect recovery of the simpler two-state clusters.
As a first comparison to alternative clustering approaches, Figure S2 shows the application of DPGMM and K-means clustering to the same well-defined two-state dataset. For the DPGMM analysis, the number of prior components was set to three with a Dirichlet concentration parameter of 0.1, such that one component would be expected to acquire negligible weight and be suppressed. This deliberate over-specification reflects a realistic experimental scenario in which the true number of states is unknown, and a robust Bayesian method should collapse redundant components rather than retain spurious ones. Instead, the DPGMM method identified three clusters as optimal. When projected onto the lifetime–intensity distribution (Fig. S2A), two inferred clusters were assigned to the same high-intensity state with nearly coincident centers, indicating the formation of a low-occupancy spurious (“ghost”) cluster. Although the cluster centers themselves align well with the underlying populations, this redundant partitioning highlights the sensitivity of DPGMM solutions to prior specification and concentration parameters, with identification of the artifact requiring explicit visualization of the clustering output. The computational cost of the DPGMM analysis was also approximately twice that of the standard GMM protocol for this simple case. For comparison, K-means clustering was applied via the elbow-point selection (Fig. S2B), which clearly identified two clusters as optimal. The resulting clustering solution (Fig. S2C) correctly separated the two populations, yielding visually reasonable partitions, albeit with a computational time again approximately twice that of our GMM-based protocol. While K-means performs adequately in this trivial, well-separated scenario, its distance-based hard assignments and restrictive cluster-shape assumptions limit its general applicability, whereas the DPGMM comparison already reveals sensitivity to priors even under idealized conditions.
Well-defined two-state system using different blinking models
Additional blinking models are presented in Figure S3, where off-state dwell times follow a power-law distribution and on-state dwell times follow power-law (Fig. S3A), exponential (Fig. S3C), or truncated power-law (Figs. S3E, G) statistics, where for the latter, a moderate power-law constant of 1.2 (Fig. S3E) and a very high value of 3.5 (Fig. S3G) were investigated. The corresponding clustering results (Figs. S3B, D, F, and H, respectively) are practically identical, each showing two clusters as the optimal choice ($\langle D^{2}\rangle=2$, FO = 0.05) and yielding perfect ground-truth recovery (ARI = 1). This suggests that variations in blinking statistics had no effect on the accuracy of state recovery for this simple two-state system.
@@
Figure 1: Results of the clustering protocol applied to a two-state simulated fluorescence lifetime–intensity dataset (10-min time, 300 particles). A. Representative intensity trace (40-ms binning shown) illustrating rapid photoblinking between emissive and background states. B. Lifetime–intensity distribution, showing two distinct populations. C. IC score plots, indicating two clusters as optimal, with $\langle D^{2}\rangle=2$ and FO =0.05=0.05. D. Optimally clustered distribution with cluster centers (black crosses) and 0.95 confidence ellipses; cluster centers at (500 cps, 0.8 ns) and (2601 cps, 2.1 ns), ARI =1=1.
Polarization and dipole orientation effects
@@
To incorporate the effect of excitation light polarization and particle orientation, we applied Eq. (4) to the higher-intensity state from Figure 1B for circularly polarized light ($\epsilon=1$) with randomized dipole orientations. For circularly polarized excitation, the excitation probability depends only on the polar angle $\theta$, following a $\sin^{2}\theta$ dependence. This introduces broadening of the population toward lower fluorescence intensities. An example intensity trace, for a particle assumed to have a fixed orientation defined by $\theta=\frac{2\pi}{3}$, $\phi=\frac{3\pi}{4}$, yielding an excitation probability of $P_{\text{{Exc}}}=0.75$ according to Eq. (4), is shown in Figure 2, where photoblinking occurs between the off-state and the reduced on-state corresponding to a simulated maximum emission intensity of 1950 cps. The smeared lifetime–intensity distribution for 300 randomly oriented particles (Fig. 2B) shows strong broadening of the unquenched state down to the background level, with an average background-subtracted intensity of $867$ cps, consistent with the expected $1/3$ of the maximum intensity.
Applying the clustering protocol to this smeared dataset yielded the IC score plots in Figure 2C, where the elbow points, corresponding to the points of greatest gradient increase and relative IC gain, are at two and three clusters for all three IC metrics. The result for two clusters (with $\langle D^{2}\rangle=2$ and FO = 0.047) is shown in Figure 2D, where the cluster of the higher state is centered offset to its mean intensity (1102 cps, 2.1 ns), and that of the lower state exactly coincides with the densely populated off-state (500 cps, 0.8 ns). Despite the density-weighted displacement of the second cluster’s intensity, ARI =1=1, indicating perfect point-level recovery per cluster.
The three-cluster solution (Fig. 2E, with $\langle D^{2}\rangle=1.987$ and FO = 0.045) splits the long-lifetime data into two clusters, with cluster 2 centered at the dense background population (500 cps, 2.1 ns) and cluster 3 at the mean of the remaining data (1650 cps, 2.1 ns). Although cluster 3 should align with 2600 cps, the lower density of points (caused by broadening) at this intensity shifts the cluster center toward the denser region at intermediate intensities. For this solution, ARI =0.8=0.8, showing that most points are correctly assigned, and the differences in $\langle D^{2}\rangle$ (1.987) and FO (0.045) relative to the two-cluster solution reflect only the broadening of the higher-intensity population. These metrics support the conclusion that two clusters (which correspond to ground truth) represent the optimal solution when considering both statistical criteria and cluster interpretability, even though the IC score for three clusters is slightly lower.
@@
Figure 2: Influence of the excitation light polarization on the simulated intensity of a molecular dipole. A. Example intensity trace (with 40 ms binning) for circularly polarized light and a dipole orientation of $\theta=\frac{2\pi}{3}$ and $\phi=\frac{3\pi}{4}$, simulated for a time of 10 min. B. Lifetime–intensity distribution of a set of 300 fixed but randomly oriented dipoles. C. Corresponding IC score plots, with two ($\langle D^{2}\rangle=2$, FO =0.047=0.047) and three clusters ($\langle D^{2}\rangle=1.987$, FO =0.045=0.045) identified as optimal choices. D. Clustering result for the two-cluster solution, with centers at (500 cps, 0.8 ns) and (1102 cps, 2.1 ns) shown as black crosses with 0.95 confidence ellipses; ARI = 1. E. Three-cluster solution, centers at (500 cps, 0.8 ns), (500 cps, 2.1 ns), and (1650 cps, 2.1 ns); ARI = 0.8, indicating some misassignment due to the broadened distribution of the higher-intensity population. F. Histogram of excitation probabilities for $10^{6}$ randomly oriented dipoles under circular ($\epsilon=1$, red), elliptical ($\epsilon=0.7$, blue), and linear ($\epsilon=0$, green) polarization.
Five-state system: effect of state overlap
To assess the impact of state overlap on cluster recoverability, the well-defined two-state system in Figure 1 was extended by three additional states at fixed intensity–lifetime coordinates: (1200 cps, 2.4 ns), (1500 cps, 1.4 ns), and (2200 cps, 0.8 ns). The original two states were kept unchanged in both mean values and SDs. All five states were assigned equal occupation probabilities to maintain constant overall data density. To isolate the effect of overlap, only the SDs of the three added states—scaled proportionally to their mean intensity and lifetime—were systematically increased, while their mean positions and probabilities were held fixed, thereby moving from well-defined to diffuse states. This controlled broadening progressively increased the spatial extent and overlap of the three additional, diffuse populations with one another and with the original two well-defined populations, allowing direct identification of the regime in which overlap degrades cluster separability and recovery. The resulting degree of overlap was quantified using the average pairwise Gaussian overlap (APGO) metric, calculated from standard Gaussian overlap integrals between all state pairs.
Figure 3 shows the effect of increasing diffuse-state widths on clustering recovery, examining the three values SD = 0.1 $\times$ mean, SD = 0.2 $\times$ mean, and SD = 0.3 $\times$ mean. For SD = 0.1 $\times$ mean (left column), APGO = 0.05, indicating minimal overlap between populations. AIC and BIC follow identical trajectories while ICL deviates only beyond five clusters (Fig. 3A), consistently identifying five clusters as optimal with $\langle D^{2}\rangle=2$ and FO = 0.05. For SD = 0.2 $\times$ mean (middle column), APGO rises to 0.24, reflecting moderate overlap. The IC curves (Fig. 3B) exhibit competing candidate elbow points at two and five clusters. The two-cluster solution yields marginally tighter clusters, with $\langle D^{2}\rangle=1.98$ and FO = 0.028, whereas the five-cluster solution is slightly less tight ($\langle D^{2}\rangle=1.94$) with a higher but still moderate FO = 0.04. Importantly, the gain in tightness obtained by merging the data into two clusters is small, while the increase in FO for the five-cluster solution reflects the deliberate separation of the three additional states rather than a loss of data coverage. In this regime, the five-cluster partition captures a larger fraction of the data structure by resolving the additional states individually, with only a modest relaxation in tightness. Consequently, despite the slightly tighter two-cluster solution, the five-cluster solution provides a more faithful representation of the full dataset, balancing cluster tightness with improved state resolution. For SD = 0.3 $\times$ mean (right column), the APGO further increases to 0.43, indicating substantial population merging. All three IC metrics again identify two and five clusters as the optimal elbow points (Fig. 3C), with corresponding auxiliary metrics of $\langle D^{2}\rangle=1.96$, FO = 0.03, and $\langle D^{2}\rangle=1.8$, FO = 0.3, respectively. In this case, the five-cluster solution shows markedly poorer compactness than before, while the two-cluster solution arises as the more statistically grounded solution, based on the cluster-quality metrics.
Figures 3 D–F show the corresponding clustered results. For SD = 0.1 $\times$ mean (Fig. 3D), five clusters are perfectly recovered with cluster centers exactly matching the underlying model, yielding ARI = 1. For SD = 0.2 $\times$ mean (Fig. 3E), the five-cluster solution retains near-complete recovery with broader confidence ellipses (ARI = 0.95), while the two-cluster solution merges multiple populations (ARI = 0.2, result not shown). For SD = 0.3 $\times$ mean (Fig. 3F), only two clusters are recovered (ARI = 0.2), with cluster centers at (500 cps, 0.8 ns) and (1920 cps, 1.65 ns), demonstrating substantial merging. These results illustrate that as states become more diffuse (characterized by larger SDs), state overlap increases, which reduces ground-truth recoverability, defining the practical breaking point of the clustering protocol, i.e., when states begin to overlap significantly with one another, clustering becomes more ambiguous.
Figure 3: Effect of state overlap by increasing state widths on five-state clustering recovery. A–C. IC score plots for each SD case: SD = 0.1 $\times$ mean (A) shows minimal overlap (APGO = 0.05), with AIC and BIC identifying five clusters ($\langle D^{2}\rangle=2$, FO = 0.05); SD = 0.2 $\times$ mean (B) exhibits competing minima at two ($\langle D^{2}\rangle=1.98$, FO = 0.028) and five clusters ($\langle D^{2}\rangle=1.94$, FO = 0.04); SD = 0.3 $\times$ mean (C) shows substantial overlap (APGO = 0.43), with two clusters selected as optimal ($\langle D^{2}\rangle=1.96$, FO = 0.03), and a formally accessible five-cluster solution showing poorer compactness ($\langle D^{2}\rangle=1.8$, FO = 0.03). D–F. Corresponding clustering results: SD = 0.1 $\times$ mean (D) recovers all five clusters perfectly (ARI = 1); SD = 0.2 $\times$ mean (E) shows near-complete recovery for five clusters (ARI = 0.95) (F) recovers only two clusters (ARI = 0.2) with centers at (500 cps, 0.8 ns) and (1920 cps, 1.65 ns).
Before proceeding to explore the effect of varying state probabilities, we tested the robustness of alternative clustering methods on the well-defined system shown in Figure 3D, where GMM with IC metrics perfectly recovered all five populations. For the DPGMM comparison, the weight concentration prior was again set to 0.1. When the number of prior components was set to six, all five clusters were recovered correctly, as shown in Figure  S4A. However, when the number of priors was increased to eight, mimicking a situation in which the true number of clusters may be unknown, the model selected seven clusters (Fig. S4B), assigning two low-occupancy “ghost” clusters in addition to the true populations, highlighting its sensitivity to prior specification. For K-means clustering, the elbow plot is shown in Figure  S4C, with gradient lines used to identify the elbow. Based on this gradient method, three or four clusters appear as plausible solutions, but the clustering results for these choices (Figs.  S4D and E) are clearly poor, failing to recover the ground-truth populations. When five clusters are explicitly specified (Fig. S4F), K-means successfully recovers all populations; however, this requires prior knowledge of the correct number of clusters, which is undesirable for an objective clustering framework. Together, these comparisons underscore both the sensitivity of DPGMM to prior settings and the limitations of K-means in objectively selecting the correct number of clusters, even when the underlying populations are well-defined.
Five-state system: effect of relative state probabilities
To investigate the influence of relative state occupancy on cluster recoverability, we used the system described in Figure 3D, with SD = 0.1 $\times$ mean and where all states have a probability of occurrence set to 0.2, as a baseline, and systematically varied the relative probabilities of the two initial states (1 and 5) with respect to those of the three additional states (2–4). The results are summarized in Figure 4 with initial/additional state probability combinations of 0.23/0.18 (left column), 0.35/0.1 (middle column), and 0.44/0.04 (right column). In these simulations, the positions and widths of all states remained fixed, so the degree of overlap was unchanged; only the relative state probabilities of the two initial and three additional states were altered. Figure 4 A–C correspond to the IC score plots for the respective cases. For the left and middle probability configurations, the IC scores consistently identify the five-cluster solution as the preferred choice, supported by near-ideal cluster tightness ($\langle D^{2}\rangle\approx 2$) and FO $\approx 0.05$, indicating that moderate changes in occupancy do not affect either the preferred cluster number or the intrinsic solution quality. The corresponding clustering results (Figs. 4D and E) show perfect recovery of all five ground-truth states, with the two initial states and the three additional states remaining clearly resolved despite their differing occupancies. In contrast, for the extreme probability imbalance in the right column (0.44/0.04), the IC analysis no longer identifies the five-cluster solution as a viable candidate. Instead, a three-cluster solution is selected as optimal, with $\langle D^{2}\rangle\approx 1.97$ and FO $\approx 0.04$, reflecting adequate compactness and data coverage. A two-cluster solution also appears as a candidate, exhibiting slightly stronger compactness ($\langle D^{2}\rangle\approx 2$) but a substantially elevated FO ($\approx 0.09$), indicating that a significant fraction of points fall outside the confidence regions, an undesirable outcome relative to the target FO $\approx 0.05$ and indicative of under-partitioning. The clustering result corresponding to the selected three-cluster solution is shown in Fig. 4F, where the two initial (well-defined) states are recovered cleanly, while the three low-occupancy states are grouped into a single composite cluster. Together, these results demonstrate that while state probabilities generally influence the statistical prominence of individual populations, clustering performance is largely insensitive to occupancy changes when geometric overlap is fixed. Only when a pronounced density imbalance is introduced, due to relative probability differences, are low-occupancy states effectively down-weighted, leading to their controlled aggregation without compromising the recovery of the dominant, well-defined states.
Figure 4: Effect of changing the relative probabilities of the two initial, high-occupancy states versus the three additional, low-occupancy states on ground-truth recovery. A–C. IC score plots for two initial-state/additional-state probability combinations, where the first value refers to the probability of the two initial states (1 and 5) and the second to the three additional states (2–4): 0.23/0.18 (A), 0.35/0.1 (B), and 0.44/0.04 (C). For the left and middle cases (A, B), the IC curves consistently identify the five-cluster solution as optimal, supported by near-ideal tightness ($\langle D^{2}\rangle\approx 2$) and FO $\approx 0.05$. In contrast, for the extreme occupancy imbalance in (C), a three-cluster solution is selected as optimal ($\langle D^{2}\rangle\approx 1.97$, FO $\approx 0.04$). D–F. Corresponding clustering results: in (D) and (E), all five ground-truth states are perfectly recovered (ARI = 1), with stable cluster centers (black crosses) and progressively reduced population of the three additional states. In (F), the two initial states are still recovered accurately, while the three low-occupancy, diffuse states are merged into a single cluster, reflecting effective down-weighting under strong probability imbalance.
Final test: effect of blinking models on a complex dataset
To assess whether blinking kinetics influence clustering performance in highly congested datasets, we generated an overlap-dominated ten-state system using different on-state dwell-time models. The dataset combines five well-defined states with five diffuse states, all assigned equal probabilities, with narrow and comparable SDs for the well-defined states and broader distributions for the additional states; polarization effects were included. Mean values and SDs are listed in Table S1, with states 1–5 corresponding to well-defined populations and states 6–10 to diffuse states. Figure S5A–C shows the results for power-law on- and off-state blinking, while panels D–F, G–I, and J–L correspond to exponential, truncated power-law, and high-exponent truncated power-law on-state models, respectively, with off-state dwell times following power-law statistics in all cases. Across all blinking models, the clustering outcome is effectively unchanged. In all cases, three clusters are recovered at similar centers (at approximately 510, 1170, and 2020 cps with corresponding lifetimes of 0.71, 2.3, and 1.45 ns, respectively), with mean cluster tightness in the range $\langle D^{2}\rangle\approx 1.80$–1.87, FO $\approx 0.03$, and ARI values of 0.11–0.12. The failure to recover the ground-truth states, therefore, does not originate from the choice of blinking kinetics but reflects an intrinsic limitation imposed by strong population overlap and distribution broadening. In this regime, clustering performance is governed by the geometric structure of the lifetime–intensity distributions rather than by the underlying temporal blinking statistics.
Broad applicability
Our clustering protocol is broadly applicable to diverse single-molecule observables beyond lifetime–intensity data, including any measurement where populations can be represented as points in a multidimensional space and exhibit spread or overlap. To demonstrate this, we simulated a spectral dataset based on peak position and full-width at half-maximum (FWHM) values for multiple states, sampling each within designated noise levels to mimic experimental variability. Figure S6A shows a scatter plot of FWHM–peak correlations, emulating a possible experimental result, while Figure S6B presents the corresponding two-dimensional ensemble-averaged spectrum. Figure S6C displays a density plot of the FWHM–peak position distribution. When applied to this dataset, the clustering protocol accurately identifies the statistically optimal number of populations. The IC score plots in Figure S6D consistently select four clusters, with $\langle D^{2}\rangle=2$ and FO = 0.06. The resulting cluster assignment, shown in Figure S6E, recovers all underlying states with ARI = 1, confirming high fidelity even under realistic noise conditions.
@@
We applied our clustering protocol to both the ungrouped and grouped-level Alexa data to demonstrate the effectiveness of the grouping technique described in Ref. (55). The Alexa dataset was chosen as a case study of rapid photoblinking, with dwell times in intensity states generally too short to be resolved visually or with a statistical change-point analysis (Fig. 5A). This was attained by deliberately not removing oxygen from the buffer solution. In addition, spin-coating of the sample introduced a random distribution of dipole orientations. We used circularly polarized excitation light to consider only the polar angle dependence of the dipole orientation, similarly to Figure 2. The resulting erratic intensity behavior and data smearing are shown in Figure 5, with the left column representing the resolved-level (ungrouped) data and the right column representing the grouped-level data. Clearly, change-point analysis of the data had difficulty resolving the actual intensity levels accurately (Fig. 5A, green trace), whereas this improved markedly after grouping the data using the built-in grouping feature of Full SMS (Fig. 5B, green trace). The lifetime–intensity distribution of the ungrouped data (Fig. 5C) was also significantly more spread out than that of the grouped data (Fig. 5D) and showed little discernible trend, making it difficult to distinguish between individual populations. In contrast, the unbiased grouping (Fig. 5D) statistically assigns similar resolved levels to a single grouped level, which effectively denoises the data and facilitates easier and more accurate clustering. Furthermore, the grouping also removed most of the short-lifetime data in the ungrouped distribution, which highlights the large fitting uncertainties when dealing with low photon budgets.
For the resolved ungrouped Alexa data, all three IC scores (Fig. 5E) indicate that three and five clusters are the most statistically prominent local minima. Considering the corresponding auxiliary metrics of $\langle D^{2}\rangle=1.81$, FO = 0.07, and $\langle D^{2}\rangle=1.71$, FO = 0.06, respectively, for a confidence interval of 0.9, the three-cluster solution is optimal. The resulting three-cluster analysis of the ungrouped data (Fig. 5G) reproduces the expected broad structure, with cluster centers corresponding to the mean values within the confidence ellipses rather than the densest populations, due to the inherent spread of the data. For the grouped data, the IC score plots decrease with increasing cluster number. Using a standard IC-based approach that selects the global minimum as the optimal solution tends to favor higher cluster numbers, which would clearly lead to overfitting in this case. Our algorithm identifies two clusters as the optimal solution, since it corresponds with $\langle D^{2}\rangle=2$ and FO = 0.09, while all other candidate solutions, including three, six, eight, and ten clusters, yield lower tightness values ($\langle D^{2}\rangle<1.94$) and smaller fractions outside (FO <0.08<0.08), suggesting over-segmentation and reduced interpretability. The result of using two clusters for the grouped data is shown in Figure 5H, which produced more coherent and visually consistent cluster centers than the three-cluster solution in Figure 5G while avoiding overfitting.
@@
Figure 5: Comparison of the performance of the clustering protocol on the ungrouped (left column) and grouped (right column) Alexa data with 52 particles. A and B. Example intensity traces over 10-s windows, with grey representing the 40-ms binned intensity data, green showing the ungrouped (A) or grouped (B) trace, and the dashed blue line indicating the background intensity level. C and D. Lifetime–intensity distributions for the corresponding datasets. E and F. IC score plots used to select the optimal number of clusters; for the ungrouped data, three clusters are preferred, whereas for the grouped data, two clusters are chosen. Clustering outcomes for the selected cluster numbers, with cluster centers indicated by black crosses and 0.9 confidence ellipses showing the spread of data points assigned to each cluster.